OpenStudy (anonymous):

A chef boiled 4 eggs and put them in a basket with 8 eggs that were not boiled. All of the eggs look the same. Randy selects an egg, keeps it, and then selects another egg. Which expression gives the probability that he selects 2 boiled eggs? A. 4/12*4/12 B. 4/12*3/11 C. 4/12*3/11 D. 4/12*3/12

OpenStudy (anonymous):

i think the answer is B

OpenStudy (whpalmer4):

I can't help but notice that B and C appear to be identical, is that intentional?

OpenStudy (anonymous):

Sorry i must typed in the wrong #.

OpenStudy (anonymous):

@whpalmer4 C. 4/12*4/11

OpenStudy (whpalmer4):

Yes, then I agree with B. You initially have a 4/12 chance of drawing a boiled egg. After that, there are 3 boiled eggs and 11 eggs in total, so the chance of drawing another one is 3/11.

OpenStudy (whpalmer4):

Though any decent chef knows how to tell a boiled egg from an uncooked one without cracking it open!

OpenStudy (anonymous):

Okay. Thx i have another question. can u check it?

OpenStudy (whpalmer4):

Sure, why not?

OpenStudy (anonymous):

A box contains 3 cherry frozen treats and 2 grape frozen treats. Maggie takes a treat from the box without looking, gives it to her brother, and then selects another treat. What is the probability that her brother gets a grape treat and she gets a cherry treat? A. 1/5 B. 6/25 C. 3/10<- D. 9/10

OpenStudy (whpalmer4):

so we are looking for the probability that she picks a grape treat, then without replacement picks a cherry treat. picking a grape treat is 2/5 because there are 2 grape treats out of 5 treats in all. after successfully picking a grape treat, there are 1 grape treat and 3 cherry treats, so the probability of randomly selecting a cherry treat is 3/(3+1) = 3/4. Multiplying the probabilities, (2*3)/(5*4) = 6/20 = 3/10.

OpenStudy (whpalmer4):

\[\frac{2}{5}*\frac{3}{4} = \frac{2*3}{5*4} = \frac{6}{20}=\frac{3}{10}\]